# IEEE Copyright Notice

### v2

@@ IV Results
IV-A Neural Network Training Setup
All models are implemented in PyTorch and four configurations for each dataset are trained: (1) Koopman with standard decoder, (2) Koopman with ABCD ($\mathcal{L}_{\text{attn-cons}}$), (3) 1D oscillator network with standard decoder, and (4) 2D oscillator network with coupled ABCD ($\mathcal{L}_{\text{attn-cons}}$, $\mathcal{L}_{\text{attn-coupling}}$). Images are downscaled to $3\times 32\times 32$ pixels and subsampled to 60 fps ($\Delta t=1/60$ s).We use a latent dimension of $k=8$ for the 1-segment dataset (four 2D oscillators) and $k=10$ for the 2-segment dataset (five 2D oscillators). We train all models for 300 epochs using AdamW with a batch size of 32 and separate learning rates for encoder, dynamics, and decoder. For ABCD models, we employ a 5-epoch warmup phase with decreased encoder and dynamic model learning rate to encourage background information to be captured by the learnable background parameters of the decoder. Training is terminated early at 150 epochs (1-segment) and 100 epochs (2-segment) for the 2D oscillator networks with the ABCD, as we observe better generalization. The implementation including the detailed architecture, hyperparameters, loss weights, and training schedules is available on GitHub [^2] Code available at: https://github.com/UThenrik/visual_oscillators_for_SCR. A video containing animations of the results presented in Sec. IV-B-IV-D is provided as supplemental media and is also available on YouTube [^3] Video available at: https://youtu.be/nbGgmY9bI-w.
@@ IV-B Attention Maps
Fig. 3 shows the attention maps for the 1-segment and 2-segment robots for Koopman and oscillator networks using the ABCD, visualizing the relative contribution of each latent dimension to the reconstruction of each pixel or to the static background. We can see that all models accurately identify the static background, with a continuous gradient at the base of the SCR from latent attention maps to background map.
@@
Figure 3: Attention maps for the 1-segment and 2-segment robots for Koopman and oscillator networks using the ABCD. Two attention maps are shown per image for the Koopman models, to achieve visual compactness.
@@
The latent attention maps focus on different regions of the robots, showing the different information of the robot they encode. By visualizing the COM of the squared attention maps for the 2D oscillators, we can also see that they are distributed along the main axis of the 1-segment and 2-segment SCR.
IV-C On-Image Oscillator Networks
@@
Figure 4: On-image 2D oscillator networks for the 1-segment and 2-segment robots (left) and latent space visualization (right). A validation state (current) is shown as well as a future state, 20 steps ahead. The oscillator networks which accurately capture the SCR dynamics, are shown with their masses, stiffness, and current actuation forces. Oscillator positions are scaled 1.5 times around their shared mean position and higher resolution images are used for visual clarity.
@@
Figure 4 shows the learned 2D oscillator networks visualized on the robot images for both current and future states (20 steps ahead). Oscillator masses are represented by marker size, pairwise stiffness by line width, and current actuation forces by arrow length and direction. Markers are colored in different shades of blue to enable identification of the same oscillators in the corresponding latent space visualization shown on the right. The relative positions of the oscillator COMs in image space match their relative positioning when latents are interpreted as 2D coordinates in latent space, as enforced by the attention coupling loss $\mathcal{L}_{\text{attn-coupling}}$. It is noted that ground stiffness and damping are not visualized to avoid visual clutter.
For the 1-segment robot, the on-image visualization clearly demonstrates the high interpretability of the coupled oscillator network: internal forces push the robot to the left in the current state, and the future state indeed shows the robot moving leftward. For the 2-segment robot, the visualization reveals more complex dynamics. In the current state, one internal force pushes strongly to the right at the upper segment while the lower oscillators experience rightward forces. Correspondingly, the future state shows the upper segment bending left while the lower segment’s curvature inverts to the right, accurately predicting the physical deformation.
@@
Notably, the learned structure for the 2-segment robot forms a chain of five oscillators, which is physically meaningful: In Cosserat rod theory, soft continuum robots are represented by an infinite chain of coupled oscillators, and the ABCD-based oscillator network correctly identifies a finite chain as an accurate reduced-order approximation. The spatial distribution of oscillators is also insightful: two oscillators at each end are spread out along the robot, while three middle oscillators cluster near the connecting segment and exhibit high mutual stiffness. This structure aligns with the physical construction, as the two-segment robot is connected by a stiff intermediate element.
@@ IV-D Single and Multi-Step Prediction
Figure 5: Multi-step reconstruction error over $0.5\text{\,}\mathrm{s}$ for 1-segment (top) and 2-segment (bottom) robots. Shaded regions indicate standard error of the mean over 50 validation trajectories. The ABCD improves multi-step reconstruction accuracy for the more complex 2-segment system for both Koopman and oscillator network dynamics.
@@
Figure 6: Multi-step reconstruction images (single step and every 6th step shown) over $0.5\text{\,}\mathrm{s}$ for 1-segment and 2-segment robots compared to ground truth for a validation trajectory. All 1-segment networks achieve highly accurate reconstructions. For the 2-segment robot, the ABCD improves long-term reconstruction accuracy for both Koopman and oscillator network dynamics.
@@
We evaluate all trained networks on single-step and multi-step prediction tasks on validation trajectories. We initialize each model with an observation and its latent velocity (computed via central differences and forward-mode AD), then autoregressively predict the next 30 steps ($0.5\text{\,}\mathrm{s}$) by repeatedly applying the learned dynamics. Prediction errors denote the dynamic reconstruction loss, averaged over 50 randomly sampled validation trajectories per model. Single-step errors represent the MSE of the first predicted frame, while multi-step errors are averaged over all 30 prediction steps.
Fig. 5 shows the reconstruction error over time and steps for both 1-segment and 2-segment robots and Fig. 6 shows the reconstruction images. For the 1-segment robot, all models achieve similar single-step and multi-step errors, apart from the 2D oscillator network with coupled ABCD which has a higher single-step prediction error and slightly higher multi-step error. Though looking at the reconstructed images in Fig. 6, all models sufficiently accurately predict the 1-segment robot dynamics. For the more complex 2-segment robot, the performance differences are more pronounced: Koopman with ABCD achieves the best multi-step reconstruction error of $9.84\times 10^{-4}$, significantly outperforming the standard Koopman ($5.66\times 10^{-3}$), while the 2D oscillator network with ABCD ($6.56\times 10^{-3}$) also outperforms the standard oscillator network ($2.27\times 10^{-2}$). The networks using the ABCD have slightly worse single-step reconstruction errors. However, we can observe sufficiently accurate single-step reconstructions in Fig. 6 and more accurate multi-step reconstructions for the networks using the ABCD, where the plain Koopman and oscillator networks diverge.
IV-E Latent Space Extrapolation
@@
Figure 7: Latent space extrapolation for 2-segment robot models. Two states 10 steps apart (left-bent and right-bent) are linearly extrapolated in latent space with factor $\alpha$. Blue markers indicate oscillator positions for attention-based models. ABCD models successfully extrapolate a bending transitions from left to right, while standard models produce non-bent, noisy states.
@@
To evaluate the learned latent representations’ ability to extrapolate observed trajectories, we perform linear interpolation and extrapolation in latent space. We select two latent states from the data set, separated by 10 time steps, representing a transition of the robot from a left-bent configuration to an approximately neutral position. For extrapolation factors $\alpha\in\{1.2,1.5,2,3\}$, we compute $\bm{z}_{\text{extra}}=\bm{z}^{(i)}+\alpha(\bm{z}^{(i+10)}-\bm{z}^{(i)})$ and decode the extrapolated latent states.
@@
Fig. 7 reveals a significant difference between models with and without the ABCD. Both Koopman with ABCD and the 2D oscillator network with coupled ABCD successfully extrapolate the bending movement, with a transition from the left-to-right bending motion, beyond a bending angle that is represented in the dataset (for $\alpha=3$). In contrast, the standard Koopman and oscillator networks without ABCD decoder fail to generate a meaningful extrapolation: their reconstructions become trapped at the near-neutral configuration and produce noisy images. This indicates that the attention mechanism not only improves interpretability and multi-step prediction accuracy for the 2-segment SCR but also learns more structured latent representations. This is expected to be highly beneficial for control.
@@ V Conclusions
This work introduced the Attention Broadcast Decoder (ABCD) for autoencoder-based latent dynamics learning from video, that is highly interpretable, accurate and computationally efficient. The ABCD generates pixel-precise attention maps for each latent dimension to enable spatial localization of learned features, and separates static background from dynamic robot motion. We demonstrated its plug-and-play integration with both Koopman operators and oscillator networks for dynamics learning of a single and double-segment soft continuum robot.
By coupling attention maps to 2D oscillator networks through an attention coupling loss, we enabled direct on-image visualization of learned dynamics, i.e., masses, stiffness, and forces, without prior knowledge of the system. For the 2-segment robot, the model autonomously discovered a chain structure consistent with Cosserat rod theory. The ABCD not only provides unprecedented interpretability but also significantly improves multi-step prediction accuracy for the complex 2-segment robot: $5.7\text{\,}\times$ error reduction for Koopman ($9.84\text{\times}{10}^{-4}$ vs. $5.66\text{\times}{10}^{-3}$) and $3.5\text{\,}\times$ for oscillator networks ($6.56\text{\times}{10}^{-3}$ vs. $2.27\text{\times}{10}^{-2}$). Both Koopman and oscillator network variants with ABCD further enable smooth latent space extrapolation where standard methods fail. Notably, only 4 and 5 2D oscillators are required for the 1- and 2-segment robot, respectively, denoting computationally efficient dynamics models.
The presented approach enables future work in several directions. Applications to control are promising, where the interpretable on-image visualizations may facilitate target state specification and trajectory planning. The learned matrices offer opportunities for parameter identification such as material specifications and simulation of structural modifications such as robot damage. The approach could be extended to different (soft) robotic systems, and to multi-camera setups, potentially learning 3D oscillator networks akin to finite element models. Additionally, transferring learned dynamics to different camera perspectives may be facilitated, as latents correspond to smoothly localized regions of the robot without encoding background information.

### v3

@@ IV Results
IV-A Neural-Network Training Setup
All models are implemented in PyTorch and four configurations for each dataset are trained: (1) Koopman with standard decoder, (2) Koopman with ABCD ($\mathcal{L}_{\text{attn-cons}}$), (3) 1D oscillator network with standard decoder, and (4) VONs ($\mathcal{L}_{\text{attn-cons}}$, $\mathcal{L}_{\text{attn-coupling}}$). Images are downscaled to $3\times 32\times 32$ pixels and subsampled to 60 fps ($\Delta t=1/60$ s). Based on an oscillator-count hyperparameter sensitivity study (see on Fig. 5 (right)), we use $n=3$ oscillators ($k=6$ latents) for the 1-segment dataset and $n=5$ oscillators ($k=10$ latents) for the 2-segment dataset. For fairness, the Koopman models use the same latent dimension $k$ as the corresponding VON setting. All models use the same encoder architecture (3-layer CNN + linear layer) and the standard decoder mirrors it with transposed convolutions. We train all models for 300 epochs using AdamW with a batch size of 32 and separate learning rates for encoder, dynamics, and decoder. For ABCD-based models, we use a short warmup phase with reduced encoder and dynamics learning rate to encourage background information to be captured by the learnable background parameters. Training is terminated early at 100 epochs for the VONs, as we observe better generalization. The implementation including the detailed architecture, hyperparameters, loss weights, and training schedules is available on GitHub [^2] Code available at: https://github.com/UThenrik/visual_oscillators_for_SCR. A video containing animations of the results presented in Sec. IV-B-IV-D is provided as supplemental media and is also available on YouTube [^3] Video available at: https://youtu.be/SMlXh-niDPs.
@@ IV-B Attention Maps
Fig. 3 shows the attention maps for the 1-segment and 2-segment robots for Koopman and oscillator networks using the ABCD, visualizing the relative contribution of each latent dimension to the reconstruction of each pixel or to the static background. We can see that all models identify the static background accurately, with VON models showing sharper seperation. A continuous gradient at the base of the SCR from latent attention maps to background map is observed.
@@
Figure 3: Attention maps for the 1-segment and 2-segment robots for Koopman and oscillator networks using the ABCD. Two attention maps are shown per image for the Koopman models for visual compactness.
@@
The latent attention maps focus on different regions of the robots, showing the different information of the robot they encode. By visualizing the COM of the squared attention maps for the 2D oscillators of the VONs, we can also see that they are distributed on and along the main axis of the 1-segment and 2-segment SCR.
IV-C Visual Oscillator Networks (VONs)
@@
Figure 4: VONs for the 1-segment and 2-segment robots (left) and latent space visualization (right) on a validation state. The VONs, which accurately capture the SCR dynamics, are shown with their masses, coupling stiffness, stiffness forces, inertial forces (in acceleration direction), and current actuation forces. Marker size, coupling line thickness, and force arrow lengths are proportional to the corresponding quantities. Coupling-to-ground stiffness and damping are learned but not visualized to avoid visual clutter (i.e., the learned oscillator system is not free-floating). Acceleration of the SCR is shown in image space by coloring the SCR edge based on image brightening and darkening from extrapolating and decoding the shown state based on latent accelerations. Oscillator positions are scaled 1.75 times around their shared mean position and higher resolution images are used for visual clarity.
@@
Figure 4 shows the learned VONs visualized on the corresponding robot images. Oscillator masses are represented by marker size, coupling stiffness by line width, and actuation, stiffness, and inertial forces by arrows. Markers are colored in different shades of blue to enable identification of the same oscillators in the corresponding latent space visualization shown on the right. The relative positions of the oscillator COMs in image space match their relative positioning when latents are viewed as 2D coordinates in latent space, as enforced by the attention coupling loss $\mathcal{L}_{\text{attn-coupling}}$. Stiffness forces act as restoring forces toward the rest configuration, while actuation often acts in the opposite direction. Inertial forces indicate the resulting acceleration while damping forces are not shown to prevent visual clutter. SCR acceleration is additionaly visualized on the image and reasonably matches the inertial force directions.
For the 1-segment robot, the actuation forces push the SCR to the lower right with stiffness forces pointing toward rest position. As the latter are stronger, the SCR’s tip is accelerated to the left, as visible by the green edge. This well matches the inertial oscillator forces. For the 2-segment robot, the stiffness forces point at return-to-rest for the different oscillators, and inertial forces predict a more complex movement with the upper part moving right and the tip moving upward. This also reasonably well matches the real image acceleration denoted by the green edge.
@@
Notably, for the 2-segment robot the learned structure forms a chain of five oscillators, which is a reasonable low-order representation. In Cosserat rod theory, soft continuum robots can be viewed as an infinite chain of coupled oscillators, and the ABCD-based oscillator network identifies a finite chain as a reduced-order approximation. The spatial distribution of oscillators is also insightful, as two oscillators at each end are spread out along the robot, while three middle oscillators cluster near the connecting segment and exhibit high mutual stiffness. This structure aligns with the physical construction, as the two-segment robot is connected by a stiff intermediate element. Like the real system, the chain elongates under actuation and contracts toward the rest state (see supplemental animation).
@@ IV-D Single and Multi-Step Prediction
Figure 5: (Left) Multi-step reconstruction error image MSE over $0.5\text{\,}\mathrm{s}$ for 1-segment (top) and 2-segment (bottom) robots. Shaded regions indicate standard error of the mean over 50 validation trajectories. (Right) Multi-step errors from a VON oscillator-count sensitivity study. The ABCD improves multi-step reconstruction accuracy for the more complex 2-segment system for both Koopman and oscillator network dynamics.
@@
Figure 6: Multi-step reconstruction images (1st, and every 6th step shown) over $0.5\text{\,}\mathrm{s}$ for 1- and 2-segment robots, compared to ground truth for a validation trajectory. All 1-segment models achieve highly accurate reconstructions. For 2 segments, the ABCD improves long-term reconstruction accuracy for both Koopman and oscillator network dynamics.
@@
We evaluate all trained networks on single-step and multi-step prediction tasks on validation trajectories. We initialize each model with an observation and its latent velocity (computed via central differences and forward-mode AD), then autoregressively predict the next 30 steps ($0.5\text{\,}\mathrm{s}$) by repeatedly applying the learned dynamics. All reported prediction errors are image MSEs between the predicted frame and the ground-truth observation.
Fig. 5 shows the reconstruction error over time and steps for both 1-segment and 2-segment robots, and Fig. 6 shows the reconstruction images. For the 1-segment robot, errors are similar across models, with multi-step MSE on the order of $5\cdot 10^{-4}$. The plain oscillator achieves $2.46\times 10^{-4}$ and the VON achieves $5.74\times 10^{-4}$. For the more complex 2-segment robot, Koopman+ABCD achieves the best multi-step MSE of $9.84\times 10^{-4}$ compared to $5.66\times 10^{-3}$ for standard Koopman, which corresponds to a $5.8\times$ improvement. The VON achieves $6.56\times 10^{-3}$ compared to $2.27\times 10^{-2}$ for the standard oscillator, which corresponds to a $3.5\times$ improvement. ABCD-based models have higher single-step image MSE, but lower long-horizon drift that improves multi-step prediction. This trend is consistent for binarized-image MSE and intersection over union (IoU), indicating that it reflects shape prediction rather than image artifacts. This suggests strong generalization of the VONs and may be aided by the dynamical consistency encouraged by the attention losses and by the inductive bias of the ABCD. The reconstruction images visualized in Fig. 6 also show how the ABCD-based models more accurate predict the 2-segment SCR shape for higher multi-step counts, and that all models predict the 1-segment SCR shape accurately.
IV-E Latent-Space Extrapolation
@@
Figure 7: Latent-space extrapolation for 2-segment robot models. Two states (left-bent and near-rest) are linearly extrapolated in latent space with factor $\alpha$. Blue markers indicate oscillator positions of the VON. ABCD models successfully extrapolate a bending transitions from left to right, while standard models produce non-bent, noisy states. For the VON, stiffness force sum is additionally shown, which acts as a restoring force toward equilibrium.
@@
We perform linear extrapolation by selecting two latent states from the data set, separated by 10 time steps, representing a transition of the robot from a left-bent configuration to an approximately neutral position. For extrapolation factors $\alpha\in\{1.2,1.5,2,3\}$, we compute $\bm{z}_{\text{extra}}=\bm{z}^{(i)}+\alpha(\bm{z}^{(i+10)}-\bm{z}^{(i)})$ and decode the extrapolated latent states.
@@
Fig. 7 reveals a significant difference between models: Both Koopman with ABCD and the VON successfully extrapolate the bending movement, with a transition from the left-to-right bending motion, beyond a bending angle that is represented in the dataset (for $\alpha=3$). The summed stiffness force of the VON remains a plausible restoring signal and increases for stronger extrapolations. In contrast, the standard Koopman and oscillator networks without ABCD decoder fail to generate a meaningful extrapolation. Their reconstructions become trapped at the near-neutral configuration and produce noisy images. This indicates that the attention mechanism not only improves interpretability and multi-step prediction accuracy for the 2-segment SCR but also learns more structured latent representations with reasonable return-to-rest stiffness forces. The physical realizability of these extrapolated states remains an open question, but the result may be relevant for future control applications and should be validated in dedicated control experiments.
@@ V Conclusions
This work introduced the Attention Broadcast Decoder (ABCD) and Visual Oscillator Networks (VONs) for autoencoder-based latent dynamics learning from video. The ABCD generates pixel-precise attention maps for each latent dimension to enable their spatial localization, and separates static background from dynamic robot motion. We demonstrated its plug-and-play integration with both Koopman operators and oscillator networks for dynamics learning of a single- and dual-segment soft continuum robot (SCR).
By coupling attention maps to VONs through an attention coupling loss, we enabled direct on-image visualization of learned dynamics, including masses, coupling stiffness, and forces, without prior knowledge of the system. For the 2-segment robot, the model autonomously discovered a chain structure similar to Cosserat rod theory. The ABCD not only provides unprecedented visual and mechanical interpretability but also significantly improves multi-step prediction accuracy for the complex 2-segment robot, achieving $5.8\text{\,}\times$ error reduction for Koopman and $3.5\text{\,}\times$ for oscillator networks. Both Koopman and VON variants with ABCD further enable smooth latent space extrapolation where standard methods fail. In our experiments, only 3 and 5 2D oscillators are required for the 1- and 2-segment robot, respectively, indicating computational efficiency.
This study has several limitations that also motivate future work. While the learned models may be well-suited for future control applications due to their structured latent dynamics and long-horizon stability, but dedicated control experiments are required. Further physical interpretability could be achieved through boundary conditions (e.g., a fixed base) or via additional constraints or regularization to better relate learned parameters to physical quantities. Strong latent-space extrapolation should be validated for physical realizability through dedicated out-of-distribution experiments or simulation. Augmenting the dataset with geometric ground truth (e.g., tip position) and evaluating under varying viewpoints would enable more direct geometric evaluation and broader generalization. Our experiments further assume largely static backgrounds. Extending to variable backgrounds is a promising direction and may be integrated into the ABCD attention-based separation of robot and background.
Beyond these limitations, promising further developments include applications to more diverse soft robots, multi-camera setups, and learning 3D oscillator networks akin to finite element models, as well as studying transfer to different camera perspectives.